# Quantifying Risks in Multi-turn Conversation with Large Language Models

### v1

@@ Abstract
Large Language Models (LLMs) can produce catastrophic responses in conversational settings that pose serious risks to public safety and security. Existing evaluations often fail to fully reveal these vulnerabilities because they rely on fixed attack prompt sequences, lack statistical guarantees, and do not scale to the vast space of multi-turn conversations. In this work, we propose QRLLM, a novel, principled Certification framework for Catastrophic risks in multi-turn Conversation for LLMs that bounds the probability of an LLM generating catastrophic responses under multi-turn conversation distributions with statistical guarantees. We model multi-turn conversations as probability distributions over query sequences, represented by a Markov process on a query graph whose edges encode semantic similarity to capture realistic conversational flow, and quantify catastrophic risks using confidence intervals.
We define several inexpensive and practical distributions—random node, graph path, and adaptive with rejection. Our results demonstrate that these distributions can reveal substantial catastrophic risks in frontier models, with certified lower bounds as high as 70% for the worst model, highlighting the urgent need for improved safety training strategies in frontier LLMs.
@@ 1 Introduction
Figure 1: Overall certification framework QRLLM(Quantifying Risks in multi-turn conversation with LLMs). Starting from a query set, we construct a graph in which edges connect semantically similar queries. On this graph, we define formal specifications as probability distributions over query sequences. For each sampled sequence, we query the LLM, use a judge model to determine whether the response is harmful, and aggregate the results to compute certification bounds on the probability of catastrophic risk.
@@
Main contributions. In this work, we present QRLLM, the first framework (shown in Figure 1) for certifying catastrophic risks in multi-turn conversations with LLMs. We are the first to formally specify the temporal safety of LLM responses in a conversational setting. We provide a general recipe for designing such specifications based on Markov processes on graph representations. We instantiate the framework with three different distributions—random node, graph path, and adaptive with rejection (Section 3), capturing a large number of realistic conversations exploitable by adversaries with fixed or adaptive attack strategies. QRLLM then certifies the target LLM by generating high-confidence bounds on the probability of catastrophic risks for a randomly sampled conversation from the distribution. Our main contributions are:
@@
  • We find a non-trivial lower bound on the probability of catastrophic risks across different frontier LLMs. We find that Claude-Sonnet-4 is the safest while Mistral-Large and DeepSeek-R1 exhibit the highest risks. We conduct case studies to identify common patterns, distractors (additional benign queries in the dialogue making refusals less likely) and context (preceding turns providing supporting information and making harmful targets clearer), that lead to catastrophic outputs.
@@ 2 Related Work
Multi-turn Attack. In contrast to single-turn attacks, which typically pose malicious questions at once with some confusion on LLMs  (Yuan et al., 2023; Wang et al., 2023; Liu et al., 2024), multi-turn jailbreaks obfuscate harmful intent by hiding it within a sequence of seemingly innocuous queries. Previous work shows this through human red-teaming  (Li et al., 2024), automated LLM attackers  (Russinovich et al., 2025; Ren et al., 2024; Yang et al., 2024), scenario-based setups  (Sun et al., 2024), query decomposition  (Zhou et al., 2024), and attacker-trained models  (Zhao & Zhang, 2025). These strategies significantly increase attack success rates compared to single-turn prompts.
Safety Evaluation of LLMs. Several datasets and benchmarks have been introduced to evaluate the safety of LLMs against harmful queries. Instruction-based benchmarks such as AdvBench  (Zou et al., 2023) and RedEval  (Bhardwaj & Poria, 2023) contain harmful or adversarial instructions that range from stereotypes, violence, to illegal activity. Generative benchmarks such as SAP (Deng et al., 2023) and AART (Radharapu et al., 2023) automatically construct adversarial prompts using models, enabling more diverse and adaptive evaluations. More recently, standardized evaluation frameworks have been proposed, targeting single-turn jailbreak robustness  (Chao et al., 2024)  (Mazeika et al., 2024) and multi-turn safety  (Yu et al., 2024; Burden et al., 2024).
@@ 3.2 Graph Distributions from Markov Process
Distributions (1) and (3) correspond to forward selection, while (2) uses backward selection. In forward selection, we specify an initial distribution $\mu$ over the starting query and a transition probability $\Pr((v_{t+1},U_{t+1})\mid(v_{t},U_{t}))$. In backward selection, we specify an endpoint distribution $\nu$ over the ending query and a backward transition rule $\Pr((v_{t},U_{t}\mid v_{t+1},U_{t+1})$. For any finite set $X$, let $\Pi(X)=\left\{\,\pi_{X}:X\to(0,1]\;\middle|\;\sum_{x\in X}\pi_{X}(x)=1\,\right\}$ denote the set of probability mass functions on $X$ that assign strictly positive probability to every element of $X$. In the following instantiations, whenever we write $\pi_{X}$, we mean an arbitrary choice of $\pi_{X}\in\Pi(X)$. We now describe the concrete instantiations of these distributions.
(1) Random node. The first query is selected according to a $\pi_{V}$, $\mu((v_{0},\{v_{0}\}))=\pi_{V}(v_{0})$. Each subsequent query is drawn from a $\pi_{V\setminus S}$:
$\Pr\big((w,T)\mid(v,S)\big)=\begin{cases}\pi_{V\setminus S}(w),&w\in V\setminus S,\;T=S\cup\{w\},\\[5.69054pt] 0,&\text{otherwise}.\end{cases}$
@@
(2.a) vanilla. The endpoint is selected from $V$ according to $\pi_{V}$, $\nu_{\text{all}}((v_{n-1},\{v_{n-1}\}))=\pi_{V}(v_{n-1}).$
(2.b) harmful target constraint. In many settings, it is advantageous to control the final query in the sequence. Biasing the endpoint steers the path toward a semantic region of interest (e.g., near the target query $q^{\star}$) while still generating coherent predecessors. The idea is that once the model has processed the earlier queries, the final query is the one where we most expect a desired behavior, so constraining it can help guide outcomes. Formally, we restrict the last query to lie in a designated target set $Q_{T}$ and define $\nu_{Q_{T}}((v_{n-1},\{v_{n-1}\}))=\pi_{Q_{T}}(v_{n-1})$.
@@
$\Pr((w,T)\mid(v,S))=\begin{cases}\pi_{N(v)\setminus S}(w),&w\in N(v)\setminus S,\;T=S\cup\{w\},\\[5.69054pt] 0,&\text{otherwise}.\end{cases}$
@@
Formally, we pick a $\pi_{N(v)}$ and define the weight on a given query $w$ by $\lambda_{v,S}(w)=\lambda_{h}\,\mathbf{1}_{\{w\in H(v,S)\}}\,\pi_{N(v)\setminus S}(w)+\lambda_{l}\,\mathbf{1}_{\{w\in n(v,S)\}}\,\pi_{N(v)\setminus S}(w)$ with weights $0<\lambda_{l}<\lambda_{h}$ are tunable parameters, where the high- and low-weight neighbor sets depending on the rejection are given by:
@@
Thus when $r_{v}=0$ the $\mathrm{prog}$ set receives higher weight (encourage progress), and when $r_{v}=1$ the $\mathrm{deprog}$ set receives higher weight. To guarantee that every query in the high-weight set has strictly larger weight than every query in the low-weight set, we require $\lambda_{h}\cdot\min_{a\in H}\;\pi_{N(v)\setminus S}(a)\;>\;\lambda_{l}\cdot\max_{b\in L}\;\pi_{N(v)\setminus S}(b)$. This condition is vacuous when either set is empty. The distribution on the first query is $\mu(v_{0},\{v_{0}\})=\pi_{V}(v_{0})$, and the normalized transition probability is
@@
Augmentation with jailbreaks. We extend the base distribution $\Pr(\gamma)$ with an independent augmentation layer. For each query $v_{t}$ in the sequence $\gamma=(v_{0},\dots,v_{n-1})$, let $B_{t}\sim\mathrm{Bernoulli}(p)$ be an independent indicator where $p$ is the probability of inserting a jailbreak at step $t$, and $g_{t}$ be a string drawn from the jailbreak prompts distribution $\mathcal{D}_{jb}$. The detailed definition of the probability mass $\Pr_{\mathcal{D}_{jb}}$ is given in Appendix A following previous work  (Chaudhary et al., 2024). We define the augmented query at step $t$ by
$\widetilde{v}_{t}=\begin{cases}g_{t}\mathbin{\|}v_{t},&B_{t}=1,\;\;g_{t}\sim\mathcal{D}_{jb},\\[2.84526pt] v_{t},&B_{t}=0,\end{cases}$
where $g_{t}\mathbin{\|}v_{t}$ denotes the concatenation of the two strings $g_{t}$ and $v_{t}$. Thus the augmented sequence is $\widetilde{\gamma}=(\widetilde{v}_{0},\widetilde{v}_{1},\dots,\widetilde{v}_{n-1})$ with probability
$\Pr(\widetilde{\gamma})\;=\;\Pr(\gamma)\,\prod_{t=0}^{n}\Pr(\widetilde{v}_{t}\mid v_{t}),\quad\text{where}\quad\Pr(\widetilde{v}_{t}\mid v_{t})=\begin{cases}p\cdot\Pr_{\mathcal{D}_{jb}}(g_{t}),&\widetilde{v}_{t}=(g_{t}\mathbin{\|}v_{t}),\\[2.84526pt] 1-p,&\widetilde{v}_{t}=(v_{t}).\end{cases}$
This definition allows jailbreak augmentation to be composed with any base sequence distribution.
@@ 4.1 Graph Construction
We construct our dataset from HarmBench (Mazeika et al., 2024), using the chemical_biological and cybercrime categories, which together comprise 68 scenarios (28 in chemical_biological and 40 in cybercrime). Each scenario is associated with a specific harmful target query $q^{\star}$. Since directly asking $q^{\star}$ is unrealistic in multi-turn conversations, for each $q^{\star}$, we expand a query set $Q$ whose elements are related to $q^{\star}$ but less harmful, in order to indirectly obtain answers of the target. We use actor-based prompts following the approach of (Ren et al., 2024): one harmful query is asked to derive a related book or person (the “actor”), and then queries are generated around this actor. To ensure diversity of the dataset, we prompt three LLMs: Gemini-2.5-Flash-Lite (Comanici et al., 2025), Deepseek-R1 (Guo et al., 2025), and Mistral-Large-2407 (Mistral AI, 2024). For each scenario, each LLM generates 10 actors, making a total of 30 actors, with 5 queries per actor. To illustrate, consider a scenario where $q^{\star}$ asks how to construct an explosive device. An example actor might be “Alfred Nobel”. Queries derived from this actor could include: “What is Alfred Nobel’s main contribution in history?”, or “What are the steps involved in his approach to this contribution?” After deduplication, we randomly sample 20 actors, yielding 100 queries per scenario.
To capture semantic relationships among queries in $Q$, we formally define an undirected graph $G=(V,E)$, where $V=Q$ and the edge set $E\subseteq V\times V$ is defined as $E=\{(u,v)\in Q\times Q\mid\ell_{\text{th}}<\mathrm{sim}(u,v)<h_{\text{th}}\}$. Here, $\mathrm{sim}(u,v)$ denotes the cosine similarity between sentence embeddings of $u$ and $v$, computed using the all-MiniLM-L6-v2 model (Reimers & Gurevych, 2019). The parameters $l_{\text{th}}$ and $h_{\text{th}}$ are lower and upper thresholds that control edge formation, ensuring that connections capture semantic relatedness while avoiding near-duplicate queries. We further define the target set $Q_{T}=\{v\in Q\mid\ell_{\text{th}}<\mathrm{sim}(v,q^{\star})<h_{\text{th}}\},$ which identifies queries that are highly similar to the harmful query $q^{\star}$ while remaining distinct, serving as endpoints for the graph-path (harmful target constraint) distribution in our certification framework.
@@ 4.2 Experimental setup
For each scenario, we consider the four specifications on distributions introduced in Section 3.2. we instantiate each $\pi_{X}$ as the uniform distribution over $X$, i.e., $\pi_{X}(v)=\tfrac{1}{|X|}\;\text{for all }v\in X.$ We apply jailbreak augmentation only to the random node distribution. To bound the probability an LLM produces catastrophic outcomes under a given specification, we apply the Clopper–Pearson method (Clopper & Pearson, 1934) to compute 95% confidence intervals for the probability of catastrophic responses. We use 50 sampled query sequences per specification. To determine whether an LLM’s response is a catastrophic outcome, we use GPT-4o (OpenAI, 2024) as a judge model, similar to prior studies (Yuan et al., 2024; Team, 2025).

### v2

# How Catastrophic is Your LLM? Certifying Risk in Conversation

@@ Abstract
Large Language Models (LLMs) can produce catastrophic responses in conversational settings that pose serious risks to public safety and security. Existing evaluations often fail to fully reveal these vulnerabilities because they rely on fixed attack prompt sequences, lack statistical guarantees, and do not scale to the vast space of multi-turn conversations. In this work, we propose C3LLM, a novel, principled statistical Certification framework for Catastrophic risks in multi-turn Conversation for LLMs that bounds the probability of an LLM generating catastrophic responses under multi-turn conversation distributions with statistical guarantees. We model multi-turn conversations as probability distributions over query sequences, represented by a Markov process on a query graph whose edges encode semantic similarity to capture realistic conversational flow, and quantify catastrophic risks using confidence intervals. We define several inexpensive and practical distributions—random node, graph path, and adaptive with rejection. Our results demonstrate that these distributions can reveal substantial catastrophic risks in frontier models, with certified lower bounds as high as 70% for the worst model, highlighting the urgent need for improved safety training strategies in frontier LLMs.
@@ 1 Introduction
Figure 1: Overall statistical certification framework C3LLM(Certification of Catastrophic risks in multi-turn Conversation for LLMs). Starting from a query set, we construct a graph in which edges connect semantically similar queries. On this graph, we define formal specifications as probability distributions over query sequences. For each sampled sequence, we query the LLM, use a judge model to determine whether the response is harmful, and aggregate the results to compute statistical certification bounds on the probability of catastrophic risk.
@@
Main contributions. In this work, we present C3LLM, the first framework (shown in Figure 1) for certifying catastrophic risks in multi-turn conversations with LLMs. We are the first to formally specify the temporal safety of LLM responses in a conversational setting. We provide a general recipe for designing such specifications based on Markov processes on graph representations. We instantiate the framework with three different distributions—random node, graph path, and adaptive with rejection (Section 3), capturing a large number of realistic conversations exploitable by adversaries with fixed or adaptive attack strategies. C3LLM then certifies the target LLM by generating high-confidence bounds on the probability of catastrophic risks for a randomly sampled conversation from the distribution. Our main contributions are:
@@
  • We find a non-trivial lower bound on the probability of catastrophic risks across different frontier LLMs. We find that Claude-Sonnet-4 and Nova Premier are the safest while Mistral-Large and DeepSeek-R1 exhibit the highest risks. We conduct case studies to identify common patterns, distractors (additional benign queries in the dialogue making refusals less likely) and context (preceding turns providing supporting information and making harmful targets clearer), that lead to catastrophic outputs.
@@ 2 Related Work
Multi-turn Attack. In contrast to single-turn attacks, which typically pose malicious questions at once with some confusion on LLMs  (Yuan et al., 2023; Wang et al., 2023; Liu et al., 2024), multi-turn jailbreaks obfuscate harmful intent by hiding it within a sequence of seemingly innocuous queries. Previous work shows this through human red-teaming  (Li et al., 2024), automated LLM attackers  (Russinovich et al., 2025; Ren et al., 2024; Yang et al., 2024), scenario-based setups  (Sun et al., 2024), query decomposition  (Zhou et al., 2024), and attacker-trained models  (Zhao and Zhang, 2025). These strategies significantly increase attack success rates compared to single-turn prompts.
Safety Evaluation of LLMs. Several datasets and benchmarks have been introduced to evaluate the safety of LLMs against harmful queries. Instruction-based benchmarks such as AdvBench  (Zou et al., 2023) and RedEval  (Bhardwaj and Poria, 2023) contain harmful or adversarial instructions that range from stereotypes, violence, to illegal activity. Generative benchmarks such as SAP (Deng et al., 2023) and AART (Radharapu et al., 2023) automatically construct adversarial prompts using models, enabling more diverse and adaptive evaluations. More recently, standardized evaluation frameworks have been proposed, targeting single-turn jailbreak robustness  (Chao et al., 2024)  (Mazeika et al., 2024) and multi-turn safety  (Yu et al., 2024; Burden et al., 2024).
@@ 3.2 Graph Distributions from Markov Process
Distributions (1) and (3) correspond to forward selection, while (2) uses backward selection. In forward selection, we specify an initial distribution $\mu$ over the starting query and a transition probability $\Pr((v_{t+1},U_{t+1})\mid(v_{t},U_{t}))$. In backward selection, we specify an endpoint distribution $\nu$ over the ending query and a backward transition rule $\Pr((v_{t},U_{t}\mid v_{t+1},U_{t+1})$. For any nonempty finite set $A\subseteq V$, we write $\pi(\cdot\mid A)$ for a probability mass function on $A$. When we write $\pi(w\mid A)$, we mean the probability assigned to $w\in A$ under this distribution. We do not fix a specific form for these distributions (they may be uniform or weighted), only that they are valid probability mass functions on the indicated sets. We now describe the concrete instantiations of these distributions.
(1) Random node. The first query is selected according to a distribution $\pi(\cdot\mid V)$ over all nodes, i.e., $\mu((v_{0},\{v_{0}\}))=\pi(v_{0}\mid V)$. Each subsequent query is drawn from a distribution over the unvisited nodes $V\setminus S$ (i.e., nodes not yet visited in the current sequence, as recorded in $S$):
$\Pr\big((w,T)\mid(v,S)\big)=\begin{cases}\pi(w\mid V\setminus S),&w\in V\setminus S,\;T=S\cup\{w\},\\[5.69054pt] 0,&\text{otherwise}.\end{cases}$
@@
(2.a) vanilla. The endpoint is selected from $V$ by $\nu_{\text{all}}((v_{n-1},\{v_{n-1}\}))=\pi(v_{n-1}\mid V).$
(2.b) harmful target constraint. In many settings, it is advantageous to control the final query in the sequence. Biasing the endpoint steers the path toward a semantic region of interest (e.g., near the target query $q^{\star}$) while still generating coherent predecessors. The idea is that once the model has processed the earlier queries, the final query is the one where we most expect a desired behavior, so constraining it can help guide outcomes. Formally, we restrict the last query to lie in a designated target set $Q_{T}$ and define $\nu_{Q_{T}}((v_{n-1},\{v_{n-1}\}))=\pi(v_{n-1}\mid Q_{T}).$
@@
$\Pr((w,T)\mid(v,S))=\begin{cases}\pi(w\mid N(v)\setminus S),&w\in N(v)\setminus S,\;T=S\cup\{w\},\\[5.69054pt] 0,&\text{otherwise}.\end{cases}$
@@
Formally, we pick a $\pi_{N(v)}$ and define the weight on a given query $w$ by $\lambda_{v,S}(w)=\lambda_{h}\,\mathbf{1}_{\{w\in H(v,S)\}}\,\pi(w\mid N(v)\setminus S)+\lambda_{l}\,\mathbf{1}_{\{w\in L(v,S)\}}\,\pi(w\mid N(v)\setminus S)$ with weights $0<\lambda_{l}<\lambda_{h}$ are tunable parameters, where the high- and low-weight neighbor sets depending on the rejection are given by:
@@
Thus when $r_{v}=0$ the $\mathrm{prog}$ set receives higher weight (encourage progress), and when $r_{v}=1$ the $\mathrm{deprog}$ set receives higher weight. To guarantee that every query in the high-weight set has strictly larger weight than every query in the low-weight set, we require $\lambda_{h}\cdot\min_{a\in H}\;\pi(a\mid N(v)\setminus S)\;>\;\lambda_{l}\cdot\max_{b\in L}\;\pi(b\mid N(v)\setminus S)$. This condition is vacuous when either set is empty. The distribution on the first query is $\mu(v_{0},\{v_{0}\})=\pi_{V}(v_{0})$, and the normalized transition probability is
@@
Augmentation layer. We extend the base distribution with an augmentation layer $\mathcal{D}_{\mathrm{aug}}(\cdot\mid v_{t})$. For each query $v_{t}$ in the sequence $\gamma=(v_{0},\dots,v_{n-1})$, this augmentation distribution $\mathcal{D}_{\mathrm{aug}}(\cdot\mid v_{t})$ depends on the current query $v_{t}$. The augmented sequence $\widetilde{\gamma}=(\widetilde{v}_{0},\dots,\widetilde{v}_{n-1})$ is obtained by sampling each augmented query independently conditional on $\gamma$:
$\widetilde{v}_{t}\sim\mathcal{D}_{\mathrm{aug}}(\cdot\mid v_{t}),\qquad t=0,\dots,n-1.$
Intuitively, this means that each query $v_{t}$ drawn from the base distribution can be replaced by an augmented version (for example, by inserting a jailbreak prefix before $v_{t}$, or by rewriting $v_{t}$). The resulting sequence has probability
$\Pr(\widetilde{\gamma})\;=\;\Pr(\gamma)\,\prod_{t=0}^{n-1}\Pr_{\mathcal{D}_{\mathrm{aug}}}(\widetilde{v}_{t}\mid v_{t}).$
This formulation covers both the identity case (when $\mathcal{D}_{\mathrm{aug}}(\cdot\mid v)$ returns $v$ with probability $1$) and stochastic modifications of $v$. We instantiate $\mathcal{D}_{\mathrm{aug}}$ using a jailbreak augmentation distribution $\mathcal{D}_{\mathrm{jb}}(\cdot\mid v)$ defined in Appendix A, which with some probability returns $v$ unchanged and with complementary probability prepends a jailbreak prompt to $v$. More generally, $\mathcal{D}_{\mathrm{aug}}$ can be defined by more structured generators, such as a second LLM that chooses mutations based on the conversation context, as long as this generator is treated as part of the attack process and induces a well-defined distribution over conversations.
@@ 4.1 Graph Construction
We construct our dataset from HarmBench (Mazeika et al., 2024), using three categories: chemical_biological, cybercrime, and illegal. Our main evaluation focuses on chemical_biological and cybercrime categories, which together comprise 68 scenarios (28 in chemical_biological and 40 in cybercrime), while results on illegal category are included as an additional domain check in Appendix C. Each scenario is associated with a specific harmful target query $q^{\star}$. Since directly asking $q^{\star}$ is unrealistic in multi-turn conversations, for each $q^{\star}$, we expand a query set $Q$ whose elements are related to $q^{\star}$ but less harmful, in order to indirectly obtain answers of the target. We use actor-based prompts following the approach of (Ren et al., 2024): one harmful query is asked to derive a related book or person (the “actor”), and then queries are generated around this actor. To ensure diversity of the dataset, we prompt three LLMs: Gemini-2.5-Flash-Lite (Comanici et al., 2025), Deepseek-R1 (Guo et al., 2025), and Mistral-Large-2407 (Mistral AI, 2024). For each scenario, each LLM generates 10 actors, making a total of 30 actors, with 5 queries per actor. To illustrate, consider a scenario where $q^{\star}$ asks how to construct an explosive device. An example actor might be “Alfred Nobel”. Queries derived from this actor could include: “What is Alfred Nobel’s main contribution in history?”, or “What are the steps involved in his approach to this contribution?” After deduplication, we randomly sample 20 actors, yielding 100 queries per scenario.
To capture semantic relationships among queries in $Q$, we formally define an undirected graph $G=(V,E)$, where $V=Q$ and the edge set $E\subseteq V\times V$ is defined as $E=\{(u,v)\in Q\times Q\mid\ell_{\text{th}}<\mathrm{sim}(u,v)<h_{\text{th}}\}$. Here, $\mathrm{sim}(u,v)$ denotes the cosine similarity between sentence embeddings of $u$ and $v$, computed using the all-MiniLM-L6-v2 model (Reimers and Gurevych, 2019). The parameters $l_{\text{th}}$ and $h_{\text{th}}$ are lower and upper thresholds that control edge formation, ensuring that connections capture semantic relatedness while avoiding near-duplicate queries. We further define the target set $Q_{T}=\{v\in Q\mid\ell_{\text{th}}<\mathrm{sim}(v,q^{\star})<h_{\text{th}}\},$ which identifies queries that are highly similar to the harmful query $q^{\star}$ while remaining distinct, serving as endpoints for the graph-path (harmful target constraint) distribution in our statistical certification framework.
@@ 4.2 Experimental setup
For each scenario, we consider the four specifications on distributions introduced in Section 3.2. We instantiate each conditional distribution $\pi(\cdot\mid X)$ as the uniform distribution over $X$, i.e., $\color[rgb]{0,0,1}\pi(v\mid X)=\tfrac{1}{|X|}\;\text{for all }v\in X.$} In our main experiments, we instantiate the augmentation layer with the jailbreak distribution $\mathcal{D}_{\mathrm{jb}}$ defined in Appendix A, using a jailbreak insertion probability $p=0.2$, and apply it only to the Random Node distribution. To bound the probability an LLM produces catastrophic outcomes under a given specification, we apply the Clopper–Pearson method (Clopper and Pearson, 1934) to compute 95% confidence intervals for the probability of catastrophic responses. We use 50 sampled query sequences per specification. To determine whether an LLM’s response is a catastrophic outcome, we use GPT-4o (OpenAI, 2024) as a judge model, similar to prior studies (Yuan et al., 2024; Team, 2025).